OpenStudy (anonymous):

wow I am so lost with this one it has fractions in it. 5/6x-1/5+4=7/10 I got 1/60. but I think it is wrong. HELP me please

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

5/6x=7/10-4+1/5 5/6x=(7-40+2)/10 5/6x=(-31)/10 -(5*10)/(6*31)=x -50/186=x x=-50/186=-25/93

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

welcome......

OpenStudy (anonymous):

so u are my fan????

OpenStudy (anonymous):

no not yet just got started on this site so i am no ones fan yet how do i do that